# On the Lebesgue Constant of Extended-Domain Spectral Methods for Elliptic PDEs

### v2

@@ 5.1 Framework in $d$-Dimensions
  • 3.
    Collocation Grid and Matrices: The grid is a tensor product of $N$ points in each dimension, resulting in $M=N^{d}$ points in $R_{0}$. The $M\times M$ matrix $\mathbf{W}_{d}$ mapping coefficients to function values is the Kronecker product of the 1D matrices: $\mathbf{W}_{d}=\mathbf{W}\otimes\dots\otimes\mathbf{W}$. For the Poisson operator $\mathcal{L}_{P}=-\Delta$, the associated basis functions are $\phi_{\mathbf{j}}=-\Delta w_{\mathbf{j}}=\lambda_{\mathbf{j}}w_{\mathbf{j}}$. The collocation matrix is thus $\mathbf{A}_{P}^{(d)}=\mathbf{W}_{d}\mathbf{\Lambda}_{d}$, where $\mathbf{\Lambda}_{d}$ is the diagonal matrix of eigenvalues $\lambda_{\mathbf{j}}$.
5.2 Asymptotic Instability and Pre-Asymptotic Behavior in d-Dimensions
The asymptotic instability proven in one dimension extends rigorously to higher dimensions. However, the pre-asymptotic stability analysis for the convection-diffusion operator faces a well-known challenge in numerical analysis for dimensions d ¿ 1.
Theorem 5.1 (Exponential Instability in d-Dimensions).
For both the $d$-dimensional Poisson and convection-diffusion operators, the extended-domain spectral collocation method is asymptotically unstable. The associated Lebesgue constant, $\Lambda_{N,d}$, grows exponentially with the number of modes in each dimension, $N$.
@@ 5.2 Asymptotic Instability and Pre-Asymptotic Behavior in d-Dimensions
The proof extends the 1D argument from Section 3 by analyzing the properties of the d-dimensional collocation matrix, which is constructed via Kronecker products.
1. Bounding the Lebesgue Constant. As established in Lemma 3.2, the Lebesgue constant is bounded below by the spectral norm of the inverse collocation matrix:
$\Lambda_{N,d}\geq C_{d}\left\lVert(\mathbf{A}^{(d)})^{-1}\right\rVert_{2},$
where $C_{d}$ is a constant that does not decay exponentially with $N$. The proof of instability therefore reduces to showing that $\left\lVert(\mathbf{A}^{(d)})^{-1}\right\rVert_{2}$ grows exponentially with $N$. For both the Poisson and convection-diffusion operators, the highest-order term is the Laplacian, so their asymptotic instability is governed by the same mechanism. We analyze the Poisson operator, $\mathbf{A}_{P}^{(d)}=\mathbf{W}_{d}\mathbf{\Lambda}_{d}$.
2. Bounding the Inverse Matrix Norm. The spectral norm of the inverse is
$\left\lVert(\mathbf{A}_{P}^{(d)})^{-1}\right\rVert_{2}=\left\lVert(\mathbf{\Lambda}_{d})^{-1}(\mathbf{W}_{d})^{-1}\right\rVert_{2}\geq\sigma_{\min}((\mathbf{\Lambda}_{d})^{-1})\left\lVert(\mathbf{W}_{d})^{-1}\right\rVert_{2}.$
The smallest singular value of the diagonal matrix $(\mathbf{\Lambda}_{d})^{-1}$ corresponds to the largest eigenvalue $\lambda_{\mathbf{j},\max}$, which scales as $\Theta(N^{2})$. Thus, $\sigma_{\min}((\mathbf{\Lambda}_{d})^{-1})=\Theta(N^{-2})$, a polynomial decay.
The crucial term is $\left\lVert(\mathbf{W}_{d})^{-1}\right\rVert_{2}$. The matrix $\mathbf{W}_{d}$ is the $d$-fold Kronecker product of the 1D matrix $\mathbf{W}$, i.e., $\mathbf{W}_{d}=\mathbf{W}\otimes\dots\otimes\mathbf{W}$. The singular values of a Kronecker product are the products of the singular values of the constituent matrices. Therefore, the smallest singular value of $\mathbf{W}_{d}$ is related to the smallest singular value of $\mathbf{W}$ by:
$\sigma_{\min}(\mathbf{W}_{d})=(\sigma_{\min}(\mathbf{W}))^{d}.$
From the 1D analysis in the proof of Theorem 3.1, we established that $\sigma_{\min}(\mathbf{W})$ decays exponentially:
$\sigma_{\min}(\mathbf{W})\leq C_{1}\sqrt{N}e^{-\alpha_{1}N}.$
Substituting this into the expression for $\sigma_{\min}(\mathbf{W}_{d})$ gives an even faster decay:
$\sigma_{\min}(\mathbf{W}_{d})\leq(C_{1}\sqrt{N})^{d}e^{-\alpha_{1}dN}.$
The spectral norm of the inverse, $\left\lVert(\mathbf{W}_{d})^{-1}\right\rVert_{2}$, is the reciprocal of the smallest singular value. Therefore, it grows exponentially:
$\left\lVert(\mathbf{W}_{d})^{-1}\right\rVert_{2}=\frac{1}{\sigma_{\min}(\mathbf{W}_{d})}\geq(C_{1}\sqrt{N})^{-d}e^{\alpha_{1}dN}.$
3. Conclusion. Combining the polynomial decay from the eigenvalues with the exponential growth from the basis matrix, we get:
$\left\lVert(\mathbf{A}_{P}^{(d)})^{-1}\right\rVert_{2}\geq\Theta(N^{-2})\cdot(C_{1}\sqrt{N})^{-d}e^{\alpha_{1}dN}.$
The exponential term $e^{\alpha_{1}dN}$ dominates all polynomial factors in $N$. Consequently, $\left\lVert(\mathbf{A}_{P}^{(d)})^{-1}\right\rVert_{2}$ and, by extension, the Lebesgue constant $\Lambda_{N,d}$, grow exponentially with $N$. This confirms the method’s asymptotic instability in any dimension $d\geq 1$. ∎
@@
Let the spectral collocation method be applied to the convection-diffusion operator $\mathcal{L}_{CD}=-\Delta+\mathbf{k}(\mathbf{x})\cdot\nabla$ in $d$ dimensions with homogeneous Dirichlet boundary conditions. If the convection field $\mathbf{k}(\mathbf{x})$ is sufficiently strong and non-vanishing, the inverse of the discrete physical-space operator, $(\mathbf{L}_{CD}^{(d)})^{-1}$, is numerically quasi-sparse with entries that decay exponentially away from the diagonal.
@@
The proof relies on applying Lemma 4.1 to the sequence of physical-space operators $\{\mathbf{L}_{CD,N}^{(d)}\}_{N\in\mathbb{N}}$. We must establish the two hypotheses of that lemma.
1. Exponential Decay of Operator Entries. The entries of the physical-space operator are given by $(\mathbf{L}_{CD,N}^{(d)})_{ij}=(\mathcal{L}_{CD}p_{\mathbf{j}})(\mathbf{x}_{\mathbf{i}})$, where $p_{\mathbf{j}}(\mathbf{x})$ are the $d$-dimensional cardinal basis functions. From Lemma 4.2 and the tensor-product construction, these functions exhibit exponential localization. Since $\mathcal{L}_{CD}$ is a local differential operator, the entries of $\mathbf{L}_{CD,N}^{(d)}$ also decay exponentially, satisfying the first hypothesis.
2. Uniform Boundedness of the Inverse. The second, and more challenging, hypothesis is the uniform boundedness of $\left\lVert(\mathbf{L}_{CD,N}^{(d)})^{-1}\right\rVert_{\infty}$. This is equivalent to proving that the scheme satisfies a discrete maximum principle. We prove this in the following lemma. ∎
@@
$(\alpha_{0}-\epsilon)-(C_{0}+C_{\epsilon})C_{P}>0,$ (9)
@@
1. Construction of the Comparison Function. We assume, without loss of generality, that the convection is strong in the $x_{1}$ direction with $k_{1}(\mathbf{x})\geq k_{\min}>0$. We use the standard barrier function $\psi(\mathbf{x})=e^{\gamma x_{1}}$ for $0<\gamma<k_{\min}$, which ensures $\mathcal{L}_{CD}\psi(\mathbf{x})\geq g_{\min}>0$. Let $M_{\infty}=\left\lVert\mathbf{f}\right\rVert_{\infty}$. We define the comparison function as
@@
3. Conclusion. By the hypothesis of the lemma, the condition in Eq. (9) holds, meaning the coefficient multiplying the norm is strictly positive. This inequality can therefore only be satisfied if $\left\lVert\nabla v_{N}^{-}\right\rVert_{L^{2}}^{2}=0$, which implies $v_{N}^{-}=0$ almost everywhere. Thus, the comparison function $v_{N}(\mathbf{x})$ is non-negative.
@@
This completes the proof. ∎
Remark 5.4 (Necessity of the Maximum Principle Approach in Higher Dimensions).
@@
It is important to clarify why the analysis for the multi-dimensional case relies on a conditional maximum principle (Lemma 5.3) rather than a direct extension of the $H^{1}$ stability argument used in 1D (Lemma 4.4). The 1D proof seamlessly transitioned from $H^{1}$ stability to uniform $\ell_{\infty}$ boundedness of the inverse operator via the one-dimensional Sobolev embedding theorem, which guarantees that $\left\lVert u\right\rVert_{L^{\infty}(I)}\leq C\left\lVert u\right\rVert_{H^{1}(I)}$.
This critical embedding, however, fails for dimensions $d\geq 2$. A function in $H^{1}(R_{0})$ is not necessarily bounded in $L^{\infty}(R_{0})$. Therefore, proving a uniform $H^{1}$ stability estimate for the numerical scheme is no longer sufficient to establish the uniform $\ell_{\infty}$ boundedness of the inverse physical-space operator, which is the key hypothesis needed for Lemma 4.1.
This necessitates a different and more powerful tool: the discrete maximum principle. By proving stability directly in the maximum norm, we bypass the failed Sobolev embedding and establish the required uniform boundedness needed to prove the exponential decay of the discrete Green’s function. The trade-off is that the proof becomes conditional on the properties of the continuous operator itself, as formalized in the hypothesis of Lemma 5.3.
Remark 5.5 (Stricter Regularity Requirements in Higher Dimensions).
The analysis of convergence rates versus coefficient regularity also extends, revealing a potential curse of dimensionality for non-smooth problems. For the solution $u$ of the convection-diffusion equation to be in the Sobolev space $H^{r+2}(R_{0})$, the product term $\mathbf{k}\cdot\nabla u$ must be in $H^{r}(R_{0})$. By the Sobolev multiplication theorem, this requires the coefficient vector field $\mathbf{k}(\mathbf{x})$ to belong to $H^{s}(R_{0})]^{d}$ for some $s>d/2$. This implies that the smoothness requirement on the PDE coefficients to ensure a certain order of convergence becomes progressively more stringent as the spatial dimension $d$ increases.
6 Numerical Validation
We now present a series of numerical experiments to validate the key findings of our theoretical framework. The tests are designed to provide computational evidence for: (1) the asymptotic exponential instability of the method for both operators, as proven in Section 3; (2) the significant pre-asymptotic difference in stability between the operators, explained by the inverse matrix structural dichotomy in Section 4; and (3) the high-order accuracy of the method for smooth solutions within this stable pre-asymptotic regime.
Our theoretical analysis focuses on the Lebesgue constant, $\Lambda_{N}$, as the definitive measure of stability. Therefore, our numerical tests compute this quantity directly to validate the theory.
6.1 One-Dimensional Problems
We first perform a set of 1D tests for several values of the extension parameter, $\delta\in\{0.5,1.0,2.0,4.0\}$. All tests in this section use equispaced collocation points.
1. Poisson’s Equation. We solve $\mathcal{L}[u]=-u^{\prime\prime}=f$ on the physical domain $I=[0,\pi]$. The exact solution is chosen to be $u(x)=\sin(x)e^{x}$.
2. Convection-Diffusion Equation. We solve $\mathcal{L}[u]=-u^{\prime\prime}+k_{0}u^{\prime}=f$ on the domain $I=[0,1]$ with a constant convection coefficient $k_{0}=10$. The exact solution is $u(x)=\sin(2\pi x)$.
Figure 1 presents the results. Panel (a) plots the Lebesgue constant $\Lambda_{N}$ versus $N$. The plot on a semi-log scale clearly shows exponential growth, validating the core conclusion of Theorem 3.1.
The plot unequivocally confirms the two central tenets of our analysis:
  • 1.
    Pre-Asymptotic Stability Difference: For any given $N$ and $\delta$, the Lebesgue constant for the convection-diffusion operator is orders of magnitude smaller than for the Poisson operator. This confirms the consequence of the structural dichotomy of the discrete Green’s function proven in Section 4.
  • 2.
    Asymptotic Instability: Both sets of curves show rapid, exponential growth. The Lebesgue constant for the Poisson operator is not only larger, but its rate of growth is visibly more severe. This is consistent with the exponential instability proven in Section 3, which ultimately affects both operators.
Panel (b) shows the corresponding $L^{2}$ errors. Despite the proven asymptotic instability, for both problems and all tested values of $\delta$, the method achieves the expected exponential convergence rates for these smooth solutions. This demonstrates that high accuracy can be obtained in the pre-asymptotic regime before the ill-conditioning becomes catastrophic in finite-precision arithmetic.
Figure 1: One-dimensional validation for various extension parameters $\delta$. (a) Lebesgue constant $\Lambda_{N}$ versus $N$. The exponential growth on the semi-log scale confirms the asymptotic instability predicted by our theory. Crucially, the Lebesgue constant for the convection-diffusion operator is orders of magnitude smaller than for the Poisson operator, validating the pre-asymptotic theory. (b) $L^{2}$ error versus NN, showing spectral convergence for both problems.
6.2 Direct Evidence of the Inverse Structure Dichotomy
The central thesis of this paper is that the observed stability difference is caused by a fundamental structural dichotomy in the inverse of the physical-space operators. While Figure 1(a) shows the consequence of this dichotomy, Figure 2 provides direct, visual evidence of the mechanism itself. We compute the physical-space operators $\mathbf{L}_{P}$ and $\mathbf{L}_{CD}$ and their inverses for $N=32$ and $\delta=1.0$.
The results are precisely as predicted by the theory:
  • 1.
    Forward Operators: Panels (a) and (c) show that both $\mathbf{L}_{P}$ and $\mathbf{L}_{CD}$ are localized, with their energy concentrated near the main diagonal. This confirms that differential operators are local in the physical-space basis, as argued in Lemma 4.5.
  • 2.
    Inverse Operators: Panels (b) and (d) reveal the dramatic difference. The inverse of the Poisson operator, $(\mathbf{L}_{P})^{-1}$, is patently dense, confirming the global nature of the elliptic Green’s function. In stark contrast, the inverse of the convection-diffusion operator, $(\mathbf{L}_{CD})^{-1}$, is a textbook example of a numerically quasi-sparse matrix. Its entries exhibit striking exponential decay away from the main diagonal.
This figure provides the ”smoking gun” evidence for Theorem 4.6, visually demonstrating that the well-behaved inverse structure of the convection-dominated operator is the source of its enhanced pre-asymptotic stability.
Figure 2: Direct visualization of the structural dichotomy in the physical-space operators and their inverses for $N=32$ and $\delta=1.0$. The colormaps show $\log_{10}$ of the absolute value of the matrix entries. (a), (c): Both forward operators are localized. (b): The inverse of the Poisson operator is dense. (d): The inverse of the convection-diffusion operator is numerically quasi-sparse, confirming the central theoretical claim.
6.3 Two-Dimensional Problems
Finally, we confirm that the method’s high-order accuracy is retained in multiple dimensions. We solve three distinct problems on the unit square $R_{0}=[0,1]\times[0,1]$ for $\delta\in\{0.5,1.0,2.0,4.0\}$. The exact solution for all tests is taken to be $u(x,y)=\sin(\pi x)\sin(\pi y)$. The test cases are:
  1. 1.
    Poisson Equation: $\Delta u=f$.
  2. 2.
    Constant-Coefficient Convection-Diffusion: $\Delta u+\mathbf{k}\cdot\nabla u=f$, with $\mathbf{k}=(10,5)$.
  3. 3.
    Varying-Coefficient Convection-Diffusion: $\Delta u+\mathbf{k}(\mathbf{x})\cdot\nabla u=f$, with $\mathbf{k}(x,y)=(5\cos(\pi x),5\sin(\pi y))$.
The $L^{2}$ convergence results are presented in Figure 3. In all three cases, the method demonstrates spectral convergence. The error curves decrease exponentially with $N$ (the number of modes per dimension), validating that the promise of high accuracy extends robustly to two-dimensional tensor-product domains. Although not plotted, the Lebesgue constants for the 2D problems exhibit the same behavior as in the 1D case: they are uniformly larger and grow even more rapidly, consistent with the analysis in Section 5, and the pre-asymptotic stability of the convection-diffusion operator remains evident.
Figure 3: $L^{2}$ convergence for 2D problems for various extension parameters $\delta$. (a) Poisson equation. (b) Convection-diffusion with constant coefficients. (c) Convection-diffusion with spatially varying coefficients. In all cases, the method achieves spectral accuracy, demonstrating its utility in the pre-asymptotic regime.
7 Conclusion

### v3

@@ 5.1. Framework in $d$-Dimensions
  • Collocation Grid and Matrices: The grid is a tensor product of $N$ points in each dimension, resulting in $M=N^{d}$ points in $R_{0}$. The $M\times M$ synthesis matrix $\bm{W}_{d}$ mapping coefficients to function values is the Kronecker product of the 1D matrices: $\bm{W}_{d}=\bm{W}\otimes\dots\otimes\bm{W}$. For the Poisson operator $\mathcal{L}_{P}=-\Delta$, the associated basis functions are $\phi_{\mathbf{j}}=-\Delta w_{\mathbf{j}}=\lambda_{\mathbf{j}}w_{\mathbf{j}}$. The collocation matrix is thus $\bm{A}_{P}^{(d)}=\bm{W}_{d}\mathbf{\Lambda}_{d}$, where $\mathbf{\Lambda}_{d}$ is the diagonal matrix of eigenvalues $\lambda_{\mathbf{j}}$.
5.2. Compounding Instability of the Basis
We first establish that the baseline instability identified in Section 3 is amplified in higher dimensions.
Theorem 5.1 (Exponential Growth of Condition Number in d-Dimensions).
For the $d$-dimensional Poisson operator, the condition number of the spectral collocation matrix $\bm{A}_{P}^{(d)}$ grows exponentially with respect to the number of modes. Specifically,
$\kappa(\bm{A}_{P}^{(d)})\geq Ce^{\alpha dN}.$
@@ 5.2. Compounding Instability of the Basis
The condition number is bounded below by the spectral norm of the inverse matrix. Using the factorization $\bm{A}_{P}^{(d)}=\bm{W}_{d}\mathbf{\Lambda}_{d}$, we have:
$\kappa(\bm{A}_{P}^{(d)})\geq\left\lVert(\bm{A}_{P}^{(d)})^{-1}\right\rVert_{2}=\left\lVert\mathbf{\Lambda}_{d}^{-1}\bm{W}_{d}^{-1}\right\rVert_{2}\geq\sigma_{\min}(\mathbf{\Lambda}_{d}^{-1})\left\lVert\bm{W}_{d}^{-1}\right\rVert_{2}.$
We analyze the two components:
  1. (1)
    The eigenvalues scale polynomially. The smallest singular value of $\mathbf{\Lambda}_{d}^{-1}$ corresponds to the inverse of the largest eigenvalue, which scales as $\Theta(N^{-2})$.
  2. (2)
    The frame matrix scales exponentially. The matrix $\bm{W}_{d}$ is the $d$-fold Kronecker product of the 1D matrix $\bm{W}$. The singular values of a Kronecker product are the products of the singular values of the constituent matrices. Thus:
    $\sigma_{\min}(\bm{W}_{d})=(\sigma_{\min}(\bm{W}))^{d}.$
    From the 1D analysis (Theorem 3.1), we know $\sigma_{\min}(\bm{W})\leq C_{1}e^{-\alpha N}$. Therefore, the smallest singular value of the $d$-dimensional frame is:
    $\sigma_{\min}(\bm{W}_{d})\leq(C_{1})^{d}e^{-\alpha dN}.$
    Consequently, $\left\lVert\bm{W}_{d}^{-1}\right\rVert_{2}=1/\sigma_{\min}(\bm{W}_{d})\geq C_{2}e^{\alpha dN}$.
Combining these, the polynomial decay of the eigenvalues is dominated by the exponential growth of the inverse frame matrix. Thus, the condition number grows as $e^{\alpha dN}$. Q.E.D.
5.3. Pre-Asymptotic Structural Dichotomy in d-Dimensions
Despite the severe baseline instability of the basis, the pre-asymptotic structural dichotomy remains valid. The inverse of the convection-diffusion operator retains its quasi-sparse structure, which explains the method’s practical robustness in multi-dimensional simulations.
@@ 5.3. Pre-Asymptotic Structural Dichotomy in d-Dimensions
Let the spectral collocation method be applied to the convection-diffusion operator $\mathcal{L}_{CD}=-\Delta+\mathbf{k}(\mathbf{x})\cdot\nabla$ in $d$ dimensions with homogeneous Dirichlet boundary conditions. If the convection field $\mathbf{k}(\mathbf{x})$ is sufficiently strong and non-vanishing, the inverse of the discrete physical-space operator, $(\bm{L}_{CD}^{(d)})^{-1}$, is numerically quasi-sparse with entries that decay exponentially away from the diagonal.
@@
The proof relies on applying Lemma 4.1 (from the 1D analysis, which holds for general matrix sequences) to the sequence of physical-space operators $\{\bm{L}_{CD,N}^{(d)}\}_{N\in\mathbb{N}}$. We must establish the two hypotheses of that lemma.
1. Exponential Decay of Operator Entries. The entries of the physical-space operator are given by $(\bm{L}_{CD,N}^{(d)})_{ij}=(\mathcal{L}_{CD}p_{\mathbf{j}})(\mathbf{x}_{\mathbf{i}})$, where $p_{\mathbf{j}}(\mathbf{x})$ are the $d$-dimensional cardinal basis functions. Since the basis is a tensor product, the cardinal functions are products of 1D cardinal functions: $p_{\mathbf{j}}(\mathbf{x})=\prod_{k=1}^{d}p_{j_{k}}(x_{k})$. From Lemma 4.2, the 1D functions decay exponentially. Therefore, the $d$-dimensional cardinal functions exhibit exponential localization. Since $\mathcal{L}_{CD}$ is a local differential operator, the entries of $\bm{L}_{CD,N}^{(d)}$ also decay exponentially, satisfying the first hypothesis.
2. Uniform Boundedness of the Inverse. The second hypothesis requires the uniform boundedness of $\left\lVert(\bm{L}_{CD,N}^{(d)})^{-1}\right\rVert_{\infty}$. This is equivalent to proving that the scheme satisfies a discrete maximum principle. Unlike the 1D case, we cannot rely on $H^{1}$ stability and Sobolev embeddings, as $H^{1}(R_{0})\not\subset L^{\infty}(R_{0})$ for $d\geq 2$. We prove the required maximum norm stability directly in the following lemma. Q.E.D.
@@
$(\alpha_{0}-\epsilon)-(C_{0}+C_{\epsilon})C_{P}>0,$ (5.1)
@@
1. Construction of the Comparison Function. We assume, without loss of generality, that the convection is strong in the $x_{1}$ direction with $k_{1}(\mathbf{x})\geq k_{\min}>0$. We use the standard barrier function $\psi(\mathbf{x})=e^{\gamma x_{1}}$ for $0<\gamma<k_{\min}$, which ensures $\mathcal{L}_{CD}\psi(\mathbf{x})\geq g_{\min}>0$. Let $M_{\infty}=\left\lVert\bm{f}\right\rVert_{\infty}$. We define the comparison function as
@@
3. Conclusion. By the hypothesis of the lemma, the condition in Eq. (5.1) holds, meaning the coefficient multiplying the norm is strictly positive. This inequality can therefore only be satisfied if $\left\lVert\nabla v_{N}^{-}\right\rVert_{L^{2}}^{2}=0$, which implies $v_{N}^{-}=0$ almost everywhere (due to the boundary conditions). Thus, the comparison function $v_{N}(\mathbf{x})$ is non-negative.
@@
This completes the proof. Q.E.D.
Remark 5.4 (Necessity of the Maximum Principle in Higher Dimensions).
@@
It is important to clarify why the analysis for the multi-dimensional case relies on a conditional maximum principle (Lemma 5.3) rather than a direct extension of the $H^{1}$ stability argument used in 1D. The 1D proof seamlessly transitioned from $H^{1}$ stability to uniform $\ell_{\infty}$ boundedness of the inverse operator via the one-dimensional Sobolev embedding theorem, which guarantees that $\left\lVert u\right\rVert_{L^{\infty}(I)}\leq C\left\lVert u\right\rVert_{H^{1}(I)}$. This critical embedding fails for dimensions $d\geq 2$. By proving stability directly in the maximum norm, we bypass the failed embedding and establish the required uniform boundedness needed to prove the exponential decay of the discrete Green’s function.
6. Numerical Validation
We now present numerical experiments to validate the theoretical framework. We analyze the conditioning of the square system ($M=N$) to verify the baseline instability, and we visualize the discrete Green’s function using a stabilized overdetermined solver to confirm the structural dichotomy.
6.1. Baseline Instability
To empirically verify the asymptotic analysis presented in Section 3 and the pre-asymptotic stabilization predicted in Section 4, we analyze the stability of the square system ($M=N$). The numerical experiments were conducted using the basis recombination method to enforce boundary conditions. It is important to note that while recombination algebraically enforces boundary values, it operates within the same underlying Fourier extension frame. Therefore, the asymptotic conditioning properties analyzed in Sections 3 and 4 remain governing factors.
While Theorem 3.1 characterizes the instability via the matrix condition number $\kappa(\bm{A}_{P})$, we verify this experimentally by computing the Lebesgue constant $\Lambda_{N}$. As established in Remark 2.4, $\Lambda_{N}$ acts as a lower bound for the algebraic instability; if $\Lambda_{N}$ grows exponentially, the condition number of the linear system must also grow exponentially. This provides a robust, solver-independent metric for the underlying frame instability.
For the Poisson problem, we set the physical domain length to $L=\pi$ and the exact solution to $u(x)=\sin(\pi x/L)e^{x}$. For the convection-diffusion problem, we set $L=1.0$ with a strong convection coefficient $k=10$, using the exact solution $u(x)=\sin(2\pi x/L)$.
The results, summarized in Figure 3, strikingly confirm the theoretical dichotomy:
  • Poisson Operator (Dashed Lines): As predicted by Theorem 3.1 (via the link in Remark 2.4), the Lebesgue constant $\Lambda_{N}$ grows exponentially with $N$ regardless of the extension parameter $\delta$. While increasing $\delta$ reduces the slope of the growth (improving the constant), it does not alter the fundamental exponential instability. Consequently, the $L^{2}$ error (Figure 3b) hits an accuracy floor (around $10^{-5}$ to $10^{-8}$ for $\delta\leq 1$) and diverges for large $N$. In contrast, the Convection-Diffusion operator delays this saturation significantly. For larger extension parameters (e.g., $\delta=4.0$), it maintains spectral convergence down to machine precision ($\approx 10^{-14}$), whereas the Poisson operator stagnates near $10^{-10}$ even for large $\delta$.
  • Convection-Diffusion Operator (Solid Lines): In stark contrast, the convection-diffusion operator with $k=10$ maintains a lower Lebesgue constant throughout the pre-asymptotic regime. This stability allows the method to achieve near-machine precision accuracy (errors $\approx 10^{-14}$) before the asymptotic frame instability eventually dominates. This validates Corollary 4.8, demonstrating that the quasi-sparse inverse structure effectively suppresses the frame’s ill-conditioning for practical discretizations.
Figure 3. Baseline instability. (a) The Lebesgue constant $\Lambda_{N}$ grows exponentially for both operators, confirming the underlying frame instability. (b) However, the Convection-Diffusion system maintains a Lebesgue constant orders of magnitude lower than Poisson in the pre-asymptotic regime.
6.2. Structural Dichotomy: The Green’s Function
To visualize the intrinsic structure of the inverse operators (Theorem 4.6) independently of the frame instability, we compute the numerical Green’s function. We utilize an overdetermined Least-Squares spectral method with explicit boundary constraints to ensure the solver itself remains robust.
Problem Setting: We solve for the response to a Gaussian source centered at $x_{c}=L/2$ on the physical domain $I=[0,\pi]$:
$\mathcal{L}[u](x)=\exp\left(-\frac{(x-x_{c})^{2}}{2\sigma^{2}}\right),\quad x_{c}=\pi/2,\quad\sigma=0.1.$ (6.1)
The computational domain is extended by $\delta=1.0$. To isolate the structural properties, we employ a stable frame with $N=60$ modes and enforce the boundary conditions $u(0)=u(\pi)=0$ via high-weight constraints in the least-squares system. We compare:
  1. (1)
    Poisson: $\mathcal{L}_{P}=-d^{2}/dx^{2}$.
  2. (2)
    Convection-Diffusion: $\mathcal{L}_{CD}=-d^{2}/dx^{2}+k(d/dx)$, with a moderate convection coefficient $k=5.0$.
Figure 4 plots the computed response $u(x)$.
  • The Poisson response (Blue) is a global, symmetric arch. In the log-scale plot (Panel b), the decay away from the peak is slow and algebraic, indicating a numerically dense inverse matrix.
  • The Convection-Diffusion response (Orange) is an asymmetric, localized spike. Panel (b) reveals a distinct linear slope on the logarithmic scale upstream of the source, confirming the exponential decay predicted by our theory. This ”numerical sparsity” effectively filters the global coupling of errors inherent to the ill-conditioned basis.
Figure 4. Visual verification of the Structural Dichotomy (Theorem 4.6). (a) Linear Scale: The Poisson response (Blue) is a global symmetric arch, indicating that the inverse matrix is dense; information at the center affects the entire domain. The Convection-Diffusion response (Orange) is a localized, asymmetric pulse. (b) Log Scale: The upstream side of the Convection-Diffusion response exhibits a straight linear slope, confirming the exponential decay predicted by our theory. This sparsity isolates numerical errors, explaining the method’s robustness in convection-dominated regimes.
6.3. 2D Validation
We extend the validation to two dimensions to confirm that the structural robustness is preserved in tensor-product domains, as discussed in Section 5. The solver utilizes a tensor-product basis formed by the 1D eigenfunctions, $w_{\mathbf{j}}(\mathbf{x})=w_{j_{1}}(x_{1})w_{j_{2}}(x_{2})$, on the domain $\Omega=[0,1]^{2}$.
We compare three test cases:
  1. (1)
    The Poisson equation $\mathcal{L}_{P}=-\Delta$.
  2. (2)
    Constant-coefficient convection-diffusion with $\mathbf{k}=(10,10)^{T}$.
  3. (3)
    Variable-coefficient convection-diffusion with $\mathbf{k}(x,y)=(5\cos(\pi x),5\cos(\pi y))^{T}$.
Figure 5 presents the $L^{2}$ convergence history. The 2D Poisson operator (Figure 5a) inherits the severe instability of the 1D case; the solver fails to reach high accuracy for small $\delta$ and diverges rapidly as $N$ increases, confirming the compounding instability described in Theorem 5.1.
Conversely, both the constant (Figure 5b) and variable (Figure 5c) coefficient convection-diffusion cases exhibit robust spectral convergence. The stabilizing effect of the convection term is even more pronounced in 2D, consistent with the theoretical requirement of Lemma 5.3, which relies on a sufficiently strong convection field to satisfy the discrete maximum principle. The variable coefficient case demonstrates that this stabilization is robust even when the convection field varies spatially, provided the local Péclet number remains sufficiently high in the pre-asymptotic regime.
Figure 5. 2D Convergence. (a) Poisson. (b) Constant Coeff CD. (c) Variable Coeff CD. The method converges spectrally, demonstrating that the structural robustness extends to tensor-product domains.
7. Conclusion